OpenStudy (anonymous):

Simplify logarithms to rewrite the following expressions in terms of u and t. Let ln x=t and ln y = u. [(ln x)^3 - ln (x^4)] / (lnx/e^2)ln(xe^2) This problems is stumping me!!!! I don't know what to do, since some lnx are inside parentheses and others aren't.

OpenStudy (zzr0ck3r):

i see no ln(y)

OpenStudy (anonymous):

sorry, the only t. i didn't mean to add ln (Y)

OpenStudy (zzr0ck3r):

ln(x)^3 - ln(x^4) = t^3 - 4ln(x) = t^3-4t this is the top bit ln(x)/e^2 = t/e^2 and ln(x*e^2) = ln(x) + ln(e^2) = t+2 so putting it all together (t^3-4t)/((t/e^2)*t+2))

OpenStudy (zzr0ck3r):

are you sure its not ln(x/e^2) on the bottom?

OpenStudy (anonymous):

no, for some reason, the ln is together with x/e^2 in parentheses!! Thanks so much!!

OpenStudy (anonymous):

The equation is more like LN and then next to it is a tiny x/e^2. It isn't lnx over e^2, but it is bunched in parentheses!

OpenStudy (zzr0ck3r):

im confused

OpenStudy (zzr0ck3r):

this does not seem to be the way a book would write it because the bottom term it is hard to tell if the ln(x*e^2) is in the num or den if you enter this in a calculator it will be on the top but i figured you mean bottom. can you put it in exactly as the book does?

OpenStudy (anonymous):